OpenStudy (anonymous):

Find dy/dx using logarithmic differentiation. y = (2 + x)^8/x, x > 0

OpenStudy (anonymous):

i guess they want you to write \[\ln(y)=\frac{8}{x}\ln(2+x)\] and then take the derivative of that one

OpenStudy (anonymous):

once you are done, you have to multiply by \((2+x)^{\frac{8}{x}}\) to get the solution to the original question